OpenStudy (anonymous):

Solve: 2x - y = -5 -2x - y = -1 I don't understand what to do?

OpenStudy (anonymous):

just add the two equations u will gt the value of y.try it

OpenStudy (anonymous):

-y = -5 - 2x y = 5 + 2x

OpenStudy (anonymous):

Plug y into the second equation.

OpenStudy (anonymous):

-2x - (5+2x) = -1

OpenStudy (anonymous):

2x+(-2x)-y-y=-5-1

OpenStudy (anonymous):

Or do what she said.

OpenStudy (anonymous):

solve dis u will get y.then put the value of y in any one eqn.u will get x.

OpenStudy (wolf1728):

Add the 2 equations A) 2x - y = -5 B) -2x - y = -1 -2y = -6 and it's easy to solve for y and for x

OpenStudy (anonymous):

Sorry, I've been working on my advanced stuff lately. So, I'm not used to approaching methods in a ' simple ' type of fashion. Lol.

OpenStudy (anonymous):

give some response samanthA

OpenStudy (anonymous):

I think she's writing the answer down.

OpenStudy (anonymous):

And she left, lol.

OpenStudy (anonymous):

ok jst try it...u will surely get the ans.

OpenStudy (anonymous):

y = 3, x = -1

OpenStudy (anonymous):

madrockz is a feather, don't listen to her.

OpenStudy (anonymous):

yes u did it.good :-)

OpenStudy (wolf1728):

Yes, well done Samantha